### v3

# An Economical Approach to Design with Precision Criteria

Abstract
Estimation frameworks for statistical inference are preferred to hypothesis testing when quantifying uncertainty and precisely estimating effect sizes are more valuable than binary decisions about statistical significance. Study design for estimation-based investigations often uses precision criteria to select sample sizes that control the length of interval estimates with respect to a sampling distribution. In this paper, we formally define a distribution that characterizes the probability of obtaining a sufficiently narrow interval estimate as a function of the sample size. This distribution can be used to determine the smallest sample size needed to ensure an interval estimate is sufficiently narrow. We prove that this distribution is approximately normal in large-sample settings for many data generation processes. However, this approximate normality may not hold for studies with moderate sample sizes, particularly when incorporating prior information or obtaining asymmetric interval estimates. Thus, we also propose an efficient simulation-based approach to approximate the distribution for the sample size by estimating the sampling distribution of interval estimate lengths at only two sample sizes. Our methodology provides a unified framework for design with precision criteria in Bayesian and frequentist settings. We illustrate the broad applicability of this framework with several examples.
Keywords: Credible intervals; confidence intervals; sample size determination; study design; the Bernstein-von Mises theorem
1 Introduction
Practitioners across disciplines design studies to collect data that meaningfully inform decision making in the presence of uncertainty. Different study objectives – such as assessing the plausibility of hypotheses or precisely estimating quantities of interest – motivate distinct approaches to study design. While many studies are designed with the goal of rejecting a null hypothesis, the null hypothesis significance testing (NHST) framework has been scrutinized for its association with the scientific replicability crisis (Open Science Collaboration, 2015). The American Statistical Association replied to these concerns by releasing an official statement on $p$-values (Wasserstein and Lazar, 2016) that challenged the statistical community to develop alternatives and extensions to the traditional hypothesis testing framework. This statement prompted various responses including a special issue of The American Statistician that discussed “moving to a world beyond p<0.05p<0.05(Wasserstein et al., 2019), a call to action from the President of the Institute of Mathematical Statistics (Meng, 2019), and a statement published in the Annals of Applied Statistics on statistical significance (Benjamini et al., 2021).
Amid the outpouring of work that followed, several contributions proposed methods for data analysis that leverage interval estimates for some scalar target of inference $\theta$ to inform decision making in Bayesian or frequentist settings (Blume et al., 2019; Goodman et al., 2019; Matthews, 2019). In both statistical paradigms, sample size determination within the estimation framework is used to ensure these interval estimates are precise enough to reliably inform decision making. In standard frequentist settings, the required sample size $n$ can be found using explicit formulas or by numerically solving equations (McHugh, 1961; Thompson, 1987; Dattalo, 2008; Riley et al., 2021; Mondal et al., 2024). Analytical formulas for the sample size may also be available in simple Bayesian settings with conjugate priors (Adcock, 1987; Joseph and Belisle, 1997). However, required sample sizes must often be found using simulation in complex Bayesian and frequentist contexts (Wang and Gelfand, 2002; De Santis and Pacifico, 2004).
Despite the differences between statistical paradigms, the asymptotic equivalence between Bayesian credible sets of credibility level $1-\alpha$ and frequentist confidence sets of confidence level $1-\alpha$ has been established (van der Vaart, 1998). However, sample size determination in the estimation framework lacks a unified approach that spans both Bayesian and frequentist settings. Regardless of the paradigm used for data analysis, the data are unknown prior to collection. To assess the precision of the interval estimate for $\theta$ at the design stage, a sampling distribution for the data is specified. In Bayesian contexts, Gubbiotti and De Santis (2011) defined two methodologies for choosing this sampling distribution: the conditional and predictive approaches. The predictive approach – but not the conditional one – accounts for uncertainty in the data-generating parameters. The conditional approach is typically used in frequentist sample size calculations, and the predictive approach is commonly viewed as more consistent with the Bayesian paradigm. However, the predictive approach is also used in frequentist contexts (Chuang-Stein, 2006; Ren and Oakley, 2014), and Bayesian studies can be designed under the conditional approach (FDA, 2019; Stevens and Hagar, 2022; Deng et al., 2024). There is thus a material need for design methodology that accommodates the conditional and predictive approaches in both statistical paradigms.
To accommodate Bayesian and frequentist approaches, we broadly introduce precision criteria for interval estimates. We henceforth use the term CI for statements that apply to both confidence intervals and credible intervals. Our methods accommodate various types of CIs – including frequentist confidence intervals as well as equal-tailed credible intervals or those defined using regions of highest posterior density. For now, we denote the length of the $100\times(1-\alpha)\%$ CI for the target of inference $\theta$ as $L_{1-\alpha}$. We later emphasize that this CI is a function of the data. Precision criteria generally select a sample size $n$ to ensure a large probability that $L_{1-\alpha}$ is sufficiently narrow with respect to the sampling distribution for the future data. The length probability criterion (LPC) was defined to formalize this objective in Bayesian settings (De Santis and Pacifico, 2004; Brutti et al., 2014). In a Bayesian or frequentist setting, the LPC could be applied to select the smallest sample size $n$ such that $\Pr(L_{1-\alpha}\leq l)\geq q$ for some target length $l$ and probability $q\in(0,1)$. This focus on controlling the length of CIs is not to the exclusion of the frequentist notion of coverage, which should approach the nominal level of $1-\alpha$ in frequentist and Bayesian contexts as $n\rightarrow\infty$ under standard regularity conditions (van der Vaart, 1998).
For Bayesian studies, Stevens and Hagar (2022) recently assigned a distribution to the LPC with respect to the sample size $n$ under the conditional approach. This distribution was called the sufficient sample size distribution (SSSD), and it was formally defined so that its $q$-quantile is the smallest sample size $n$ such that $L_{1-\alpha}$ will have length at most $l$ with probability at least $q$ (i.e., such that $\Pr(L_{1-\alpha}\leq l)\geq q$). The SSSD was observed to be approximately normal in a variety of situations via simulation (Stevens and Hagar, 2022), but theoretical justification for this approximate normality was not explored. Furthermore, conditions under which the SSSD exists and this approximate normality holds were not clearly defined. In this work, we define the length probability distribution (abbreviated as the LP distribution) that extends the SSSD to accommodate Bayesian and frequentist CIs under both the conditional and predictive approaches to specify sampling distributions.
By estimating the LP distribution and its quantiles, practitioners can select sample sizes for scientific studies that attain precision criteria. To support flexible study design based on precision criteria, sample sizes that satisfy the LPC can be found using intensive simulation (Wang and Gelfand, 2002). In general, many samples of size $n$ are simulated to estimate the sampling distribution of CI lengths, and the proportion of samples for which the CI length is at most $l$ estimates the quantile of the LP distribution for that sample size. That process is repeated for each $n$ considered until the desired quantile of the LP distribution is estimated. Although simulation-based procedures can be generally applied to design studies with various statistical models and CI types, such naive procedures are computationally prohibitive and wastefully ignore information provided by the large-sample theory that pertains to interval estimates (van der Vaart, 1998). It would therefore be ideal to develop a framework for design with precision criteria that leverages CI theory to efficiently guide flexible simulation-based procedures. Such an economical framework to determine the minimum sample size $n$ that satisfies the LPC would meaningfully expedite design with precision criteria in Bayesian and frequentist contexts.
In this paper, we propose two methods to estimate the LP distribution across statistical paradigms. The first method analytically approximates the LP distribution using asymptotic theory. The second method leverages this asymptotic theory to inform an efficient method to estimate the LP distribution based on simulation results conducted at only two sample sizes; this method is generally applicable and flexibly accommodates moderate sample size settings where asymptotic approximations may be unsuitable. Similar approaches have recently been proposed to efficiently select sample sizes within the hypothesis testing framework (Hagar and Stevens, 2024, 2025a, 2025c, 2025b). While our design framework proposed here is theoretically intricate, its implementation is straightforward, promoting an economical and broadly useful approach to simulation-based sample size determination with precision criteria.
The remainder of this article is structured as follows. We introduce background information and notation in Section 2. In Section 3, we state the conditions under which the LP distribution is approximately normal and prove this result as a theorem. In Section 4, we develop a theoretical proxy to the sampling distribution of CI lengths, and we propose a method in Section 5 that adapts this theory to determine the minimum sample size $n$ that satisfies the LPC. Our unified approaches to estimate the LP distribution can be utilized in Bayesian and frequentist settings under the conditional or predictive approach. In Section 6, we conduct numerical studies for several examples to assess the accuracy with which our proposed methods estimate the LP distribution and recommend sample sizes. Section 7 concludes with a summary and discussion of extensions to this work.
2 Preliminaries
Our design framework denotes data from a random, future sample as $\boldsymbol{W}^{{}_{(n)}}$, where the sample size $n$ is the number of independent sampling units (ISUs). The observed data are denoted by $\boldsymbol{w}^{{}_{(n)}}$. For non-clustered data, each observation is an ISU and in settings with clustered data, multiple dependent observations may comprise a single ISU. For each ISU, the sample $\boldsymbol{W}^{{}_{(n)}}$ may consist of response(s) $\boldsymbol{Y}$ along with a vector or matrix of additional covariates $\boldsymbol{X}$ that may encode information about treatment group assignment and/or assignment to ISUs. We assume that each ISU in $\boldsymbol{W}^{{}_{(n)}}$ is generated independently according to the model $f(\boldsymbol{w};\boldsymbol{\eta})$, where $\boldsymbol{\eta}$ denotes a vector of parameters. The target of inference $\theta$ is specified as a function $g(\cdot)$ of these parameters: $\theta=g(\boldsymbol{\eta})$. In an estimation framework, we aim to draw inference using a two-sided $100\times(1-\alpha)\%$ CI for $\theta$. We henceforth denote the length of the $100\times(1-\alpha)\%$ CI as $L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}})$ for a future sample of $n$ ISUs; the length of the corresponding CI for an observed sample is $L_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}})$. This notation now emphasizes that the CI length and its sampling distribution are functions of the data.
Algorithm 1 details a simulation-based procedure to estimate the sampling distribution of CI lengths at a given sample size $n$. We now discuss several design inputs for this algorithm. We characterize data generation for $\boldsymbol{W}^{{}_{(n)}}$ using the model $f^{+}(\boldsymbol{w};\boldsymbol{\eta}^{+})$. We have that $\boldsymbol{\eta}\subseteq\boldsymbol{\eta}^{+}$ since additional parameters beyond those needed to specify $\theta$ may be required to generate data. For instance, regression settings require additional parameters because those used to generate the covariates $\boldsymbol{X}$ are not part of the regression model. Parameters to specify the ISU sizes according to some probability distribution could also be considered. The model $f^{+}(\boldsymbol{w};\boldsymbol{\eta}^{+})$ generalizes $f(\boldsymbol{w};\boldsymbol{\eta})$ to take these additional parameters. The prior $p(\boldsymbol{\eta})$ is used to induce the posterior of $\theta=g(\boldsymbol{\eta})$ in Bayesian settings. This input is not relevant for inference based on frequentist confidence intervals. We must also specify the number of simulation repetitions $m$ for Algorithm 1.
The final design input that we must specify is $\Psi$, a probability model that characterizes how $\boldsymbol{\eta}^{+}$ values are drawn in each simulation repetition $r=1,\dots,m$. The model $\Psi$ defines scenarios with parameter values that practitioners hope to observe (De Santis, 2007; Berry et al., 2011). It differs from the prior $p(\boldsymbol{\eta})$ in Bayesian settings that synthesizes all available pre-study knowledge about the data generation process. The model $\Psi$ is respectively degenerate and nondegenerate under the conditional and predictive approaches (Gubbiotti and De Santis, 2011). Under the conditional approach, the values $\{\boldsymbol{\eta}^{+}_{r}\}_{r=1}^{m}$ generated across all $m$ simulation repetitions in Line 3 of Algorithm 1 are therefore identical. Line 3 also generates a sample $\boldsymbol{w}^{{}_{(n)}}_{r}$ in each simulation repetition given the value for $\boldsymbol{\eta}^{+}_{r}$ drawn from $\Psi$.
Algorithm 1 Sampling Distribution Estimation
[…]
2:  for $r$ in 1:$m$ do
[…]
4:   Compute estimate $\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}}_{r})$
5:  return $\{\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}}_{r})\}_{r=1}^{m}$
The collection of estimates $\{\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}}_{r})\}_{r=1}^{m}$ in Line 5 of Algorithm 1 is used to estimate the sampling distribution of CI lengths under $\Psi$. The relevant $100\times(1-\alpha)\%$ CIs for $\theta$ can be estimated using any Bayesian or frequentist method that is based on likelihood functions. This flexibility accommodates the use of simple or complex interval estimates – such as approximate frequentist confidence intervals obtained using normal theory and sample summary statistics or Bayesian highest posterior density intervals based on computationally approximating the posterior distribution of $\theta$. The probability of satisfying the length criterion can be estimated as
$\dfrac{1}{m}\sum_{r=1}^{m}\mathbb{I}\left\{\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}}_{r})\leq l\right\}.$ (1)
To determine whether the LPC is satisfied, we introduce the notation $\xi(a,b)$ to denote the $a^{\text{th}}$ order statistic of the collection of observations $b$. For a sample size $n$, the estimated probability of satisfying the length criterion in (1) is at least $q$ if and only if $\xi_{q}=\xi(\lceil mq\rceil,\{\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}}_{r})\}_{r=1}^{m})\leq l$. As $n\rightarrow\infty$, the sampling distribution of CI lengths converges to a point mass at 0 under standard regularity conditions (van der Vaart, 1998). The probability of satisfying the length criterion, defined given fixed $l$ and $\alpha$, therefore approaches 1 as the sample size $n$ increases. The optimal sample size $n$ is the smallest integer such that the estimate in (1) is at least $q$ (and $\xi_{q}\leq l$).
The optimal choice for $n$ is informed by exploring the sampling distribution of CI lengths to estimate the probability of satisfying the length criterion in (1). However, the procedure in Algorithm 1 requires independent implementation for each $n$ value we consider. This process can be computationally intensive, but we could reduce the computational burden by repurposing the results from Algorithm 1 for previously considered sample sizes to estimate (1) for new $n$ values. This process would allow us to explore new sample sizes without conducting additional simulations. We propose such a method that combines simulation and large-sample theory to design studies with precision criteria in this paper and begin its development in Section 4. But first, we formally define the LP distribution and prove that it is approximately normal under certain conditions in Section 3, which allows us to contrast the performance of our economical simulation-based method with results based solely on large-sample theory.
3 Approximate Normality of the Length Probability Distribution
The LP distribution is defined such that when the sample size $n$ is taken as its $q$-quantile, the resulting $100\times(1-\alpha)\%$ CI will have length at most $l$ with probability at least $q$. We let the LP curve be the cumulative distribution function (CDF) of the LP distribution. More precisely, let $F_{l}(n)$ be the LP curve for fixed $l$ and $1-\alpha$ such that
$F^{-1}_{l}(q):=\text{inf}\{n\in\mathbb{Z}^{+}:\Pr(L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}})\leq l)\geq q\}.$ (2)
[…]
In this section, we provide theoretical justification for the approximate normality of the LP distribution under the conditional approach where the model $\Psi$ is degenerate. This normality is useful because it facilitates fast analytical sample size calculations that we improve upon with the broader methodology described in Sections 4 and 5. We later describe how this approximate normality can also be exploited under the predictive approach. To ensure the approximate normality of the LP curve in Bayesian and frequentist settings, the regularity conditions for the asymptotic normality of the maximum likelihood estimator (MLE) of $\boldsymbol{\eta}$ must be satisfied (van der Vaart, 1998). Those conditions are detailed in Appendix A.1 of the online supplement. For Bayesian inference, we also require the conditions for the Bernstein-von Mises (BvM) theorem (van der Vaart, 1998) detailed in Appendix A.2. Both sets of conditions involve independent and identically distributed assumptions. In this paper, such assumptions must be satisfied at the level of the ISU but not at the observation level.
For a particular credibility or confidence level $1-\alpha$, the BvM theorem establishes the asymptotic equivalence between Bayesian credible sets and frequentist confidence sets. We leverage this asymptotic equivalence to prove the approximate normality of the LP distribution. The BvM theorem also motivates the limiting form for the posterior of $\theta$ in simulation repetition $r$ that we present in Lemma 1, which is written generally to encompass the conditional and predictive approaches.
Lemma 1.
Assume the conditions in Appendix A.2 are satisfied for $\boldsymbol{\eta}^{+}_{r}\sim\Psi$. Let $\hat{\theta}^{{}_{(n)}}_{r}=g(\hat{\boldsymbol{\eta}}^{{}_{(n)}}_{r})$ be the maximum likelihood estimate for $\theta=g(\boldsymbol{\eta})$ expressed in terms of the number of ISUs $n$. For large $n$, the posterior of $\theta$ is approximately
$\mathcal{N}\left(\hat{\theta}^{{}_{(n)}}_{r},n^{-1}\Lambda(\boldsymbol{\eta}_{r})\right),$ (3)
$\text{where}~\Lambda(\boldsymbol{\eta}_{r})=\frac{\partial g}{\partial\boldsymbol{\eta}}^{\emph{T}}\mathbb{E}_{\boldsymbol{X}}\left[\mathcal{I}_{\emph{ISU}}(\boldsymbol{\eta}\mid{\boldsymbol{x}})\right]^{-1}\frac{\partial g}{\partial\boldsymbol{\eta}}$
and $\mathcal{I}_{\emph{ISU}}(\boldsymbol{\eta}\mid{\boldsymbol{x}})$ is the Fisher information of an ISU with covariates $\boldsymbol{x}$}.
In Appendix B.1 of the supplement, we prove Lemma 1. We also justify that Lemma 1 holds true when the number of observations in each ISU, encoded in the covariates $\boldsymbol{X}$}, is randomly drawn according to the model $\Psi$. The constant $\Lambda(\boldsymbol{\eta}_{r})$ from Lemma 1 can be obtained analytically when there are no additional covariates, but it may need to be estimated using Monte Carlo integration in more complex settings. We use this constant for theoretical purposes here. The theory in Appendix B.1 additionally demonstrates that the approximate sampling distribution of the MLE $\hat{\theta}^{{}_{(n)}}~|~\boldsymbol{\eta}=\boldsymbol{\eta}_{r}$ is $\mathcal{N}(g(\boldsymbol{\eta}_{r}),n^{-1}\Lambda(\boldsymbol{\eta}_{r}))$ under the regularity conditions in van der Vaart (1998).
In large-sample Bayesian and frequentist settings, the length of the $100\times(1-\alpha)\%$ CI is determined by the scalar asymptotic variance $n^{-1}\Lambda(\boldsymbol{\eta}_{r})$. The model $\Psi$ is degenerate under the conditional approach, so we emphasize that $\boldsymbol{\eta}_{r}$ is a constant in the following derivation. Given $\alpha$ and large $n$,
$L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}})\approx\dfrac{2z_{1-\alpha/2}}{\sqrt{n}}\times{\Lambda(\hat{\boldsymbol{\eta}}_{r}^{{}_{(n)}})^{\nicefrac{{1}}{{2}}}},$ (4)
where $\hat{\boldsymbol{\eta}}_{r}^{{}_{(n)}}$ is the MLE for $\boldsymbol{\eta}_{r}$ and $z_{1-\alpha/2}$ is the upper $\alpha/2$ quantile of $\mathcal{N}(0,1)$. When the conditions in Appendix A.1 hold, the delta method ensures that
$\sqrt{n}\left(2z_{1-\alpha/2}\Lambda(\hat{\boldsymbol{\eta}}_{r}^{{}_{(n)}})^{\nicefrac{{1}}{{2}}}-2z_{1-\alpha/2}\Lambda(\boldsymbol{\eta}_{r})^{\nicefrac{{1}}{{2}}}\right)\xrightarrow{d}\mathcal{N}\left(0,4z^{2}_{1-\alpha/2}\dfrac{\partial}{\partial\boldsymbol{\eta}}\Lambda(\boldsymbol{\eta})^{\nicefrac{{1}}{{2}}}\bigg|_{\boldsymbol{\eta}=\boldsymbol{\eta}_{r}}^{\emph{T}}\mathbb{E}_{\boldsymbol{X}}\left[\mathcal{I}_{\text{ISU}}(\boldsymbol{\eta}\mid{\boldsymbol{x}})\right]^{-1}\dfrac{\partial}{\partial\boldsymbol{\eta}}\Lambda(\boldsymbol{\eta})^{\nicefrac{{1}}{{2}}}\bigg|_{\boldsymbol{\eta}=\boldsymbol{\eta}_{r}}\right).$ (5)
For simplicity, we denote the variance in (5) as $4z^{2}_{1-\alpha/2}\tau^{2}$ so that $\tau^{2}$ is the asymptotic variance of ${\sqrt{n}}(\lambda(\hat{\boldsymbol{\eta}}_{r}^{{}_{(n)}})^{\nicefrac{{1}}{{2}}}-\Lambda(\boldsymbol{\eta}_{r})^{\nicefrac{{1}}{{2}}})$.
The results in (4) and (5) imply that for large sample sizes $n$,
$L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}})~\dot{\sim}~\mathcal{N}\left(\dfrac{2z_{1-\alpha/2}\Lambda(\boldsymbol{\eta}_{r})^{\nicefrac{{1}}{{2}}}}{\sqrt{n}},\dfrac{4z^{2}_{1-\alpha/2}\tau^{2}}{n^{2}}\right).$ (6)
Theorem 1 uses the approximate sampling distribution for $L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}})$ in (6) to prove that the LP distribution is approximately normal under the conditional approach. This approximate normality holds when the interval length $l$ decreases to 0. As $l\rightarrow 0^{+}$, the sample sizes $n$ that satisfy the LPC increase to $\infty$.
Theorem 1.
[…]
$\lim_{l\rightarrow 0^{+}}\sup_{n\in\mathbb{Z}^{+}}\left\lvert F_{l}(n)-\Phi\left(\dfrac{n-\mu_{l}}{\sigma_{l}}\right)\right\rvert\xrightarrow{P}0.$
We prove Theorem 1 in Appendix B.2 of the supplement, but we consider its practical implications here. Theorem 1 imposes constraints on $\Lambda(\boldsymbol{\eta}_{r})$ and $\tau^{2}$ to ensure the normal distribution that approximates the LP distribution for a given length $l$ has finite mean $\mu_{l}>0$ and standard deviation $\sigma_{l}>0$. For large sample sizes $n$ corresponding to small target lengths $l$, the values for $\mu_{l}$ and $\sigma_{l}$ from Theorem 1 can be used to determine quantiles of the LP distribution under the conditional approach as stated in Corollary 1.
Corollary 1.
Let $\mu_{l}$ and $\sigma_{l}$ be defined as in Theorem 1. If we obtain a sample of size $n=\mu_{l}+z_{q}\sigma_{l}$ according to a conditional model $\Psi$, the probability that the $100\times(1-\alpha)\%$ CI for $\theta$ satisfies the length criterion approaches $q$ as $l\rightarrow 0^{+}$.
Corollary 1 can be extended for use with the predictive approach by approximating the LP distribution as a mixture of normal distributions. In particular, one can obtain $m$ parameter values $\{\boldsymbol{\eta}^{+}_{r}\}_{r=1}^{m}\sim\Psi$. Theorem 1 can be applied to obtain $\mu_{l,r}$ and $\sigma_{l,r}$ for each repetition $r$. When $m$ is sufficiently large, the LP distribution can then be estimated as the mixture of $\{\mathcal{N}(\mu_{l,r},\sigma^{2}_{l,r})\}_{r=1}^{m}$ distributions. Note that similarly to how asymptotic normality is exploited for finite samples, we do not require that $l=0$ to invoke Theorem 1. However, it is not trivial to determine how small $l$ must be to yield a sample size $n$ that is large enough to apply the asymptotic result in Theorem 1 for a given study. Corollary 1 is therefore most useful in the limiting case with large sample sizes. Nevertheless, the theoretical results from this section are valuable. In Section 4, we leverage insights from this section to develop a method that prompts better sample size recommendations in moderate sample size scenarios – particularly in settings with asymmetric CIs and Bayesian contexts with informative priors.
4 A Proxy to the Sampling Distribution of CI Lengths
Flexible, simulation-based design methods that control the length of CIs require that we estimate the sampling distribution of CI lengths for various sample sizes $n$. We approximate this sampling distribution by generating data $\boldsymbol{w}^{{}_{(n)}}$ using the straightforward process in Algorithm 1. For theoretical development, we create a proxy for the sampling distribution of the logarithm of the CI lengths. These proxies substantiate the theory that underpins our proposed methodology, but they are not used in our design methods proposed in Section 5.
The delta method and the result in (6) give rise to an approximate sampling distribution for $\kappa(\boldsymbol{W}^{{}_{(n)}})=\log(L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}}))$. For simplicity, the notation $\kappa(\boldsymbol{W}^{{}_{(n)}})$ does not incorporate the confidence or credibility level $1-\alpha$. The approximate sampling distribution for $\kappa(\boldsymbol{W}^{{}_{(n)}})$ is
$\kappa(\boldsymbol{W}^{{}_{(n)}})~\dot{\sim}~\mathcal{N}\left(\log(2z_{1-\alpha/2}{\lambda(\boldsymbol{\eta}_{r})^{\nicefrac{{1}}{{2}}}})-\dfrac{1}{2}{\log(n)},\dfrac{\tau^{2}}{n\lambda(\boldsymbol{\eta}_{r})}\right).$ (7)
For theoretical purposes, we could simulate from this approximate sampling distribution of $\kappa(\boldsymbol{W}^{{}_{(n)}})~|~\boldsymbol{\eta}^{+}=\boldsymbol{\eta}^{+}_{r}$ using CDF inversion with a point $u_{r}\in[0,1]$:
$\kappa^{{}_{(n)}}_{r}=\log(2z_{1-\alpha/2}{\lambda(\boldsymbol{\eta}_{r})^{\nicefrac{{1}}{{2}}}})-\dfrac{1}{2}{\log(n)}+\dfrac{z_{u_{r}}\tau\Lambda(\boldsymbol{\eta}_{r})^{-\nicefrac{{1}}{{2}}}}{\sqrt{n}}.$ (8)
We emphasize that $\kappa^{{}_{(n)}}_{r}$ in (8) depends on the point $u_{r}$ and parameter value $\boldsymbol{\eta}^{+}_{r}$ through the subscript $r$. The $\{\kappa^{{}_{(n)}}_{r}\}_{r=1}^{m}$ values corresponding to points $\{u_{r}\}_{r=1}^{m}\sim\mathcal{U}[0,1]$ and parameter values $\{\boldsymbol{\eta}^{+}_{r}\}_{r=1}^{m}\sim\Psi$ define our large-sample proxy to the sampling distribution of the logarithm of CI lengths. Under the predictive approach, there are two sources of randomness in the proxy sampling distribution of $\kappa^{{}_{(n)}}$. The first source is associated with the parameter values $\boldsymbol{\eta}^{+}_{r}$ for simulation repetition $r$. The second source is related to the point $u_{r}$ used to generate from the limiting distribution in (7).
For theoretical purposes, we consider how this proxy sampling distribution changes as a function of the sample size. We first consider the median $\kappa^{{}_{(n)}}_{*}$ of the proxy sampling distribution conditional on the parameter value $\boldsymbol{\eta}_{r}\subseteq\boldsymbol{\eta}^{+}_{r}$. Based on (8), this median is a linear function of $\log(n)$:
$\kappa^{{}_{(n)}}_{*}=\log(2z_{1-\alpha/2}{\lambda(\boldsymbol{\eta}_{r})^{\nicefrac{{1}}{{2}}}})-\dfrac{1}{2}{\log(n)}.$ (9)
We note that $\kappa^{{}_{(n)}}_{*}$ in (9) is not a stochastic quantity when conditioning on $\boldsymbol{\eta}_{r}^{+}$. We also define the logarithm of the absolute difference between each $\kappa^{{}_{(n)}}_{r}$ value and $\kappa^{{}_{(n)}}_{*}$ as
$\log(\lvert\kappa^{{}_{(n)}}_{r}-\kappa^{{}_{(n)}}_{*}\rvert)=\log(\lvert z_{u_{r}}\rvert\tau\Lambda(\boldsymbol{\eta}_{r})^{-\nicefrac{{1}}{{2}}})-\dfrac{1}{2}{\log(n)}.$ (10)
[…]
Corollary 2 follows directly from the results in (9) and (10). It provides guidance concerning how to economically assess whether the length criterion is satisfied at a broad range of sample sizes. In particular, Corollary 2 formally establishes that our proxy sampling distribution can be decomposed into various deterministic, linear functions of $\log(n)$; hence, exploration of the $n$-space can be achieved by estimating the sampling distribution of CI lengths at only two values of $n$.
Corollary 2.
[…]
5 Economical Estimation of the Length Probability Distribution
The methods we propose to plan studies with precision criteria can be broadly applied when the conditions in Appendix A are satisfied. We generalize the results from Theorem 1 to develop methodology in Algorithm 2 that is easily implemented and performs well with moderate to large sample sizes. Algorithm 2 allows users to efficiently explore the sample size space to find the minimum $n$ that satisfies the LPC. We obtain this sample size recommendation by estimating the LP distribution from the LP curve. Our approach involves estimating the sampling distributions of CI lengths at only two sample sizes: $n_{0}$ and $n_{1}$. Algorithm 2 details a general application of our methodology with the conditional approach, and we later describe potential modifications to accommodate the predictive approach.
Algorithm 2 Sample Size Determination with the LP Curve
[…]
2:  Let $n_{0}$ be $\lceil\mu_{l}\rceil$ corresponding to the median of $\{\Lambda(\boldsymbol{\eta}_{r})\}_{r=1}^{m}\sim\Psi$
3:  Estimate $\{\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n_{0})}}_{r})\}_{r=1}^{m}$ via Algorithm 1 and their logarithms $\{\hat{\kappa}^{{}_{(n_{0})}}_{r}\}_{r=1}^{m}$
4:  Model the median of the sampling distribution of $\kappa(\boldsymbol{W}^{{}_{(n)}})$ as $\hat{\kappa}_{*}^{{}_{(n)}}=\hat{\kappa}_{*}^{{}_{(n_{0})}}+0.5\log(n_{0})-0.5\log(n)$,    where $\hat{\kappa}_{*}^{{}_{(n_{0})}}$ is the median of $\{\hat{\kappa}^{{}_{(n_{0})}}_{r}\}_{r=1}^{m}$
5:  for $r$ in 11:$m$ do
[…]
7:  Use binary search to find $n_{1}$, the smallest $n$ such that $\xi$($\lceil mq\rceil$, $\{\hat{\kappa}^{{}_{(n)}}_{r}\}_{r=1}^{m}$) $\leq\log(l)$
8:  Estimate $\{\widehat{L}_{1-\alpha}(\boldsymbol{w}^{{}_{(n_{1})}}_{r})\}_{r=1}^{m}$ via Algorithm 1 and their logarithms $\{\hat{\kappa}^{{}_{(n_{1})}}_{r}\}_{r=1}^{m}$
9:  Model the median of the sampling distribution of $\kappa(\boldsymbol{W}^{{}_{(n)}})$ such that $\kappa_{*}^{{}_{(n)}}$ is the linear function of    $\log(n)$ passing through $(\log(n_{0}),\hat{\kappa}_{*}^{{}_{(n_{0})}})$ and $(\log(n_{1}),\hat{\kappa}_{*}^{{}_{(n_{1})}})$
10:  for $r$ in 11:$m$ do
[…]
12:  Use the functions in Lines 9 to 11 to approximate the sampling distribution of $\kappa(\boldsymbol{W}^{{}_{(n)}})$ and the LP    curve across the $n$-space
13:  Use binary search to find $n_{2}$, the smallest $n$ such that $\xi$($\lceil mq\rceil$, $\{\hat{\kappa}^{{}_{(n)}}_{r}\}_{r=1}^{m}$) $\leq\log(l)$
14:  return LP curve and $n_{2}$ as the recommended $n$
We elaborate on several steps in Algorithm 2 below. We choose the initial sample size $n_{0}$ in Line 2 as $\lceil\mu_{l}\rceil$ from Theorem 1 for the median value of $\Lambda(\boldsymbol{\eta}_{r})$ induced by $\boldsymbol{\eta}_{r}^{+}\sim\Psi$. Regardless of whether the conditional or predictive approach is used, this sample size provides a suitable starting point for LP curve estimation since $n_{0}$ should correspond to a quantile of the corresponding LP distribution such that $q$ is not too close to 0 or 1. We emphasize that our methodology is flexible and can readily be integrated with any computational or analytical method used to compute CI lengths in Line 3. The notation $\hat{\kappa}^{{}_{(n)}}_{r}$ is also introduced in Line 3. These logarithms from the true sampling distribution of $\kappa(\boldsymbol{W}^{{}_{(n)}})$ leverage independently generated samples $\boldsymbol{w}^{{}_{(n)}}_{r}$ for each simulation repetition $r$. Unlike for $\kappa^{{}_{(n)}}_{r}$ from the proxy sampling distribution in Section 4, there is no relationship between the $\hat{\kappa}^{{}_{(n)}}_{r}$ values corresponding to two different sample sizes that happen to have the same index $r$.
[…]
In Lines 9 to 11 of Algorithm 2, we model the sampling distribution of CI lengths across the $n$-space using linear approximations that are less reliant on large-sample results. These approximations use independent estimates of the sampling distribution at the sample sizes $n_{0}$ and $n_{1}$, and they exploit the linear trends in the proxy sampling distribution discussed in Section 4. For the true sampling distributions, we sort the logarithms of the absolute differences between $\hat{\kappa}^{{}_{(n)}}_{r}$ and $\hat{\kappa}^{{}_{(n)}}_{*}$ estimated at $n_{0}$ and $n_{1}$ and construct linear approximations using the same order statistic at both sample sizes. This approach is suitable when the true value of $\Lambda(\boldsymbol{\eta}_{r})$ is similar for all $\boldsymbol{\eta}^{+}_{r}\sim\Psi$. When $\Psi$ is nondegenerate under the predictive approach, the process in Lines 9 to 11 can be modified: we instead split the logarithms of the absolute differences for each sample size into subgroups based on the order statistics of their $\Lambda(\boldsymbol{\eta}_{r})$ values before constructing the linear approximations.
In Line 13 of Algorithm 2, we repeat the process in Line 7 with our improved linear approximations to obtain the final sample size recommendation $n_{2}$. These linear approximations yield unbiased estimates of the LP curve at sample sizes $n_{0}$ and $n_{1}$. Thus, the suitability of LP curve estimation and the corresponding sample size recommendation $n_{2}$ only relies on the accuracy of the empirically estimated slopes. We compare the performance of Algorithm 2 with the purely analytical method to estimate the LP distribution from Section 3 when considering study design for several examples in Section 6.
6 Illustrative Examples
6.1 Measurement Method Comparison
We first assess how well our methodology estimates the LP distribution in the context of measurement method comparison. This example allows us to consider repeated measurements and the impact of ISU sizes. In this subsection, we plan to compare two measurement methods that measure total prostate-specific antigen (tPSA), a tumour marker used to diagnose and screen for prostate cancer. This example is based on a real study described in Ferraro et al. (2023) that compares reference and new measurement methods.
We suppose that the following linear mixed-effects structural model characterizes the data and relates measurements by the two methods:
$\begin{split}\log(Y_{i1k})&=S_{i}+M_{i1k}\\ \log(Y_{i2k})&=\gamma+\beta S_{i}+M_{i2k},\end{split}$ (11)
where $i=1,2,\dots,n$ indexes the participants, $j=1,2$ respectively correspond to the reference and new measurement method, and $k=1,2,\dots,K$ indexes the replicate measurements. Thus, $Y_{ijk}$ is a random variable that represents the future value of method $j$’s $k^{\text{th}}$ measurement of participant $i$. We use a logarithmic transformation in model (11) because distributions for tPSA data are typically right skewed. $S_{i}$ is a random variable that represents the unknown true value of the measurand for subject $i$ on the logarithmic scale. In this example, we make a common assumption that $S_{i}\sim\mathcal{N}(\mu,\sigma^{2}_{s})$ and subjects are sampled randomly from the target population. $M_{ijk}$ is a random variable that represents the measurement error of method $j=1,2$. We assume that the $M_{ijk}$ are independent of each other and of $S_{i}$, and that they follow a $\mathcal{N}(0,\sigma^{2}_{j})$ distribution, where $\sigma_{j}$ quantifies the measurement variation of method $j$.
We note that the mean and variance parameters associated with the model (11) are distinct from $\mu_{l}$ and $\sigma_{l}$, the parameters that characterize the normal approximation to the LP distribution from Theorem 1. The parameters $\gamma$ and $\beta$ respectively characterize the fixed and proportional bias of the new measurement method relative to the reference one. Using our notation from Section 2, the “covariates” $\boldsymbol{X}$ for this example encode the ISU sizes (i.e., $2K$ observations for each participant across the two measurement methods). We assume the ISU sizes are the same for all participants, which is reasonable for measurement method comparison. Ultimately, our parameters for model (11) are $\boldsymbol{\eta}=(\gamma,\beta,\mu,\sigma_{s},\sigma_{1},\sigma_{2})$.
We summarize similarity between the two measurement methods in the frequentist paradigm using the probability of agreement (PoA) (Stevens et al., 2017). The PoA used as the target of inference for our study is
$\theta=\Pr(\lvert\log(Y_{i2})-\log(Y_{i1})\rvert\leq c)=\Phi\left(\dfrac{c-\gamma-(\beta-1)\mu}{\sqrt{(\beta-1)^{2}\sigma^{2}_{s}+\sigma^{2}_{1}+\sigma^{2}_{2}}}\right)-\Phi\left(\dfrac{-c-\gamma-(\beta-1)\mu}{\sqrt{(\beta-1)^{2}\sigma^{2}_{s}+\sigma^{2}_{1}+\sigma^{2}_{2}}}\right)=g(\boldsymbol{\eta}).$ (12)
The interval $(e^{-c},e^{c})$ represents a clinically acceptable relative difference between measurements on the untransformed scale. Our study therefore aims to estimate the probability that the two measurement methods yield clinically similar measurements. We emphasize that given particular values for $\boldsymbol{\eta}$, the probability in (12) is a constant that depends on the underlying level of agreement between the two measurement methods. We aim to precisely estimate this PoA in a precision-based study.
To design this study, we define $\Psi$ using the model in (11) with $\boldsymbol{\eta}_{r}=(0,1.01,2.32,1.20,0.04,0.10)$ for all simulation repetitions $r$ under the conditional approach. We define the PoA using a value of $c=\log(1.2)$. This choice for $c$ implies that a relative difference of up to 20% between the untransformed measurements is clinically acceptable. The parameter values $\boldsymbol{\eta}_{r}=\boldsymbol{\eta}$ were selected by adjusting the estimated parameters from Ferraro et al. (2023) so that the true PoA is $\theta=0.9$. We consider a two-sided $100(1-\alpha)\%=95\%$ Wald-based confidence interval for the PoA. To ensure the bounds of the CI are between 0 and 1, the CI is first constructed on the logit scale and its endpoints are transformed. Details for constructing such CIs via the delta method are provided in Stevens et al. (2017). We plan our study with the goal of obtaining an interval estimate for $\theta$ having length at most $l=0.05$ with probability $q=0.8$. To assess the impact of the ISU sizes, we consider $K=\{2,3,4,5\}$.
We estimate the LP distribution using various methods for each of the scenarios described above. We first estimate the LP distribution analytically using the results from Theorem 1. The estimated distributions are $\mathcal{N}(155.68,21.57)$ for $K=2$, $\mathcal{N}(100.01,14.18)$ for $K=3$, $\mathcal{N}(73.74,10.58)$ for $K=4$, and $\mathcal{N}(58.43,8.46)$ for $K=5$. We then approximated the LP curves using Algorithm 2. Finally, we approximated the LP curve in each scenario by naively simulating the sampling distribution of CI lengths for all $n\in(\mu_{l}-3\sigma_{l},\mu_{l}+3\sigma_{l})$ that were multiples of 5, where $\mu_{l}$ and $\sigma_{l}$ are the parameters from Theorem 1. All sampling distributions were estimated with $m=10^{4}$ simulation repetitions. Figure 1 visualizes the approximated LP curves for the three estimation methods we considered.
Figure 1: LP curves obtained using various estimation methods when the PoA is the target of inference. The horizontal dotted line represents $q=0.8$.
For all scenarios in Figure 1, the results from Algorithm 2 align better with those obtained using naive simulation than the analytical approximation to the LP curve resulting from Theorem 1. Even so, the sample size recommendations based on $q=0.8$ differ by at most $n=3$ across the three estimation methods in each of the four scenarios. Thus, the analytical approximation to the LP curve is reasonable for this example, which is not surprising given that the sample sizes are not too small and the CIs are relatively symmetric. Upon calculating the partial derivatives for the delta method, it took a fraction of a second to obtain the analytical approximation to each LP curve. Roughly 25 seconds on a high-computing server with 72 cores were required approximate the LP curve via Algorithm 2 for each of the scenarios. It took between 2 and 6 minutes using the same computing resources to approximate the LP curves in Figure 1 with naive simulation. Algorithm 2 is more efficient because we need only estimate the sampling distribution of CI lengths at two values of $n$. To consider sample sizes in increments of 5, we required between 10 and 27 sampling distribution estimates for each curve corresponding to naive simulation.
The sample size recommendation for $K=2$ based on Algorithm 2 is $n=176$. Of course, the recommended sample size decreases as the ISU size $2K$ increases (i.e., the recommended $n$ is only 68 when $K=5$). Practitioners can select an optimal design for this study by weighing the costs of recruiting a new participant against those associated with obtaining more repeated measurements. Although not explored in this example, our method also accommodates designs where the ISU size is allowed to vary according to a probability distribution specified via the model $\Psi$. Because larger ISU sizes generally lead to smaller recommended $n$, it is better to choose a model $\Psi$ that errs on the side of understating (instead of overstating) the ISU size when there is uncertainty in the size of the ISU. We explore the performance of our methods to approximate the LP curve with further numerical studies in Appendix C of the supplement. In frequentist settings, these studies confirm the strong performance of Algorithm 2, but the quality of the analytical LP curve approximation based on Theorem 1 deteriorates when CIs are based on test statistics with skewed sampling distributions.
6.2 Estimation of Illness Duration
We next assess the performance of our proposed methods in the context of estimating the duration of Mpox illness. The Mpox virus is predominantly spread via direct contact with the active lesions of an already infected individual (Kim et al., 2023). In previous clinical trials (Klein, 2024), Mpox illness duration was defined as the time in days between starting tecovirimat (a safe drug approved for use with smallpox) and experiencing the resolution of all Mpox lesions. For this example, we want to estimate the 0.9-quantile of this distribution for Mpox illness duration. Our subsequent numerical results serve purely illustrative purposes, but we emphasize the value in precisely estimating quantities associated with illness durations to inform public health recommendations. This example also allows us to consider settings with independent observations, account for additional covariates, and accommodate censoring.
The observed data for a sample of $n$ participants is $\{y_{i},\delta_{i},x_{i}\}_{i=1}^{n}$. The observed time for participant $i$ is denoted by $y_{i}$. We also define a binary variable to indicate whether patient $i$’s time is a lesion resolution time ($\delta_{i}=1$) or a right censoring time ($\delta_{i}=0$). The additional covariate $x_{i}$ is the logarithm of patient $i$’s Ct value, where these data are centered after taking the logarithm. Higher Ct values correspond to lower Mpox viral loads, and we account for this covariate because we anticipate that higher viral loads will be associated with longer illness durations. Using our notation from Section 2, the binary indicator $\delta_{i}$ could also be viewed as an additional covariate.
For illustration, we suppose that the resolution time $T_{i}$ and censoring time $C_{i}$ are exponentially distributed such that the exponential rates are log-linear functions of $x_{i}$:
$T_{i}\sim\text{EXP}(\gamma+\beta x_{i})~~~\text{and}~~~C_{i}\sim\text{EXP}(\xi+\rho x_{i}),$ (13)
where $\gamma$, $\beta$, $\xi$, and $\rho$ are regression parameters. The to-be-observed time $Y_{i}$ for participant $i$ is the minimum of $T_{i}$ and $C_{i}$. Because the censoring process in (13) depends only on the observed covariate $x_{i}$, these censored data are considered to be missing at random. However, the Bayesian methods we use in this subsection could in theory accommodate more complex missingness mechanisms. At the design stage, we must also define a distribution for $\{X_{i}\}_{i=1}^{n}$. We suppose that $X\sim\mathcal{N}(0,\sigma_{x}^{2})$ for this example, and that the data from participants $i$ and $i^{\prime}$ are independent for $i\neq i^{\prime}$. To summarize, the parameters for our analysis model are $\boldsymbol{\eta}=(\gamma,\beta)$, but the parameters $\boldsymbol{\eta}^{+}=(\gamma,\beta,\xi,\rho,\sigma_{x})$ are used for data generation.
Our target of inference used for study design is the 0.9-quantile of the distribution for the resolution time $T$ when $x=0$: $\theta=g(\boldsymbol{\eta})=\log(10)\exp(-\mu)$. We aim to estimate a $100\times(1-\alpha)\%=99\%$ posterior highest density interval (HDI) for $\theta$ having length at most $l=5$ days with probability $q=0.8$. To demonstrate the broad applicability of our methods, we design this study under the conditional and predictive approaches. For the conditional approach, we use the parameter values $\boldsymbol{\eta}^{+}_{r}=(\log(\log(10)/18),0.5,-\log(32),1,0.25)$ informed by findings from Kim et al. (2023) for all simulation repetitions $r$. These parameter values are such that the true value of $\theta_{r}$ is 18 days and lesion resolution times are censored for roughly 20% of participants. Under the predictive approach, we assume that $\gamma_{r}\sim\mathcal{N}(\log(\log(10)/18),0.07^{2})$, $\beta_{r}\sim\mathcal{N}(0.5,0.07^{2})$, $\xi_{r}\sim\mathcal{N}(-\log(32),0.07^{2})$, and $\rho_{r}\sim\mathcal{N}(1,0.07^{2})$; we note that small changes to these parameter values give rise to substantial changes in $\theta_{r}$ and the censoring proportions. We again assume that $\sigma_{x,r}=0.25$ for all repetitions under the predictive approach.
To facilitate Bayesian analysis, we consider a diffuse and informative prior for $\gamma$ to explore how our methods perform. The diffuse prior for $\gamma$ is $\mathcal{N}(0,10^{2})$, and the informative prior is $\mathcal{N}(-2,0.3^{2})$. This informative prior is relatively well specified for illustration. For all simulations in this subsection, we assign a diffuse $\mathcal{N}(0,10^{2})$ prior to $\beta$. These choices result in four approach and prior combinations. For each combination, we analytically approximated the LP curve using the results from Theorem 1. For the predictive approach, this analytical approximation arises from a mixture of normal distributions that we obtained by applying Theorem 1 to 1000 simulated $\boldsymbol{\eta}^{+}_{r}$ values. We also approximated the LP curve using Algorithm 2 with $m=10^{4}$ simulation repetitions; to accommodate the predictive approach, the CI lengths were split into 10 bins based on the true $\Lambda(\boldsymbol{\eta}_{r})$ value at each sample size before constructing the linear approximations. We lastly approximated the LP curve by simulating the sampling distribution of CI lengths in increments of $n=10$. The resulting LP curve approximations are visualized in Figure 2.
Figure 2: LP curves obtained using various estimation methods when the 0.9-quantile of Mpox illness duration is the target of inference. The horizontal dotted line represents $q=0.8$.
Algorithm 2 again yields more accurate LP curve approximation than the analytical approximation. The performance of the analytical approximation resulting from Theorem 1 is reasonable for the settings with the diffuse priors for $\gamma$. Because the analytical approximation to the LP curve does not account for the prior distribution, it overestimates the required sample size when the relatively well-specified informative prior for $\gamma$ is used ($n=478$ vs. 456 for the conditional approach and $n=506$ vs. 484 for the predictive approach). Overall, the LP distributions under the predictive approach exhibit greater dispersion than those under the conditional approach. That is, we require less (more) data to attain the length criterion with smaller (larger) probabilities $q$. This greater dispersion in the LP distribution is a direct result of there being greater dispersion in the sampling distribution of CI lengths under the predictive approach.
It took less than one second to approximate the LP distribution under the conditional approach using Theorem 1. We needed to numerically compute 1000 $\lambda(\boldsymbol{\eta}_{r})$ values and their corresponding derivatives in $\tau^{2}$ to approximate the LP distribution under the predictive approach, which took 15 seconds on a high computing server with 72 cores. Roughly one hour on this server was required to approximate all LP curves in Figure 2 via Algorithm 2. This process was computationally intensive because we approximated each posterior of $\boldsymbol{\eta}$ using Markov chain Monte Carlo with 1000 burnin iterations and 5000 retained draws, but it could have been expedited using analytical methods for posterior approximation. Nevertheless, Algorithm 2 is still much more efficient than naive simulation. On the same computing server, it took 9 and 19 hours to naively approximate the LP curve for this example under the conditional and predictive approaches, respectively. Our numerical results in Appendix C confirm the strong performance of Algorithm 2 in Bayesian settings and further demonstrate that analytical approximations based on Theorem 1 may perform poorly with moderate sample sizes when (i) informative priors are used or (ii) the posterior distributions used to construct the CI exhibit substantial skewness.
7 Discussion
In this paper, we developed an economical framework for study design with precision criteria. This framework determines optimal sample sizes that ensure interval estimates are narrow enough to reliably inform decision making. This framework is founded on the approximate normality of the LP distribution. We propose two methods to estimate this distribution and recommend sample sizes. One method that performs well for large-sample studies is predicated purely on asymptotic theory, and the other flexibly and efficiently estimates the LP distribution using simulations conducted at only two sample sizes. The efficiency of our simulation-based method stems from considering a proxy for the sampling distribution of Bayesian and frequentist CI lengths based on large-sample theory. Our numerical studies illustrate that our simulation-based method to estimate the LP distribution performs well in moderate sample size settings – even when considering informative priors or asymmetric CIs.
Our proposed framework can be broadly applied to design studies with precision criteria when CIs are constructed using a method that is based on likelihood functions. Future work could extend this methodology to accommodate design with CIs that are based on the frequentist nonparametric or the Bayesian bootstrap. Our approach could also be extended in other aspects to accommodate more complex design criteria. For instance, it would be useful to consider economical sample size determination methods for studies that aim to satisfy criteria for both the study power and the precision of interval estimates.
Additionally, we could further develop design methods with precision criteria in a sequential framework or in a framework with multiple targets of inference. Such extensions would require additional theoretical development because we would need to account for the level of dependence in the joint sampling distribution of CI lengths across all potential analyses and targets of inference. Finally, our design framework requires that the regularity conditions for the asymptotic normality of the MLE and BvM theorem are satisfied. It would therefore be of interest to consider design methods with precision criteria in settings where these regularity conditions are relaxed.
Supplementary Material
These materials include a detailed description of the conditions for Theorem 1 along with its proof and additional simulation results. The code to conduct the numerical studies in the paper is available online: https://github.com/lmhagar/LPDistribution.
Funding Acknowledgement
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) by way of a PGS D scholarship, postdoctoral fellowship, and Grant RGPIN-2025-04137.

### v4

@@
 Abstract
-Estimation frameworks for statistical inference are preferred to hypothesis testing when quantifying uncertainty and precisely estimating effect sizes are more valuable than binary decisions about statistical significance.
+Estimation frameworks for statistical inference are preferred to hypothesis testing when quantifying uncertainty and precise estimation are more valuable than binary decisions about statistical significance.
 Study design for estimation-based investigations often uses precision criteria to select sample sizes that control the length of interval estimates with respect to a sampling distribution.
@@
 The American Statistical Association replied to these concerns by releasing an official statement on $p$-values (Wasserstein and Lazar, 2016) that challenged the statistical community to develop alternatives and extensions to the traditional hypothesis testing framework.
-This statement prompted various responses including a special issue of The American Statistician that discussed “moving to a world beyond p<0.05p<0.05(Wasserstein et al., 2019), a call to action from the President of the Institute of Mathematical Statistics (Meng, 2019), and a statement published in the Annals of Applied Statistics on statistical significance (Benjamini et al., 2021).
+This statement prompted various responses including a special issue of The American Statistician that discussed “moving to a world beyond p<0.05p<0.05(Wasserstein et al., 2019) along with calls to action from the President of the Institute of Mathematical Statistics (Meng, 2019) and from the Royal Statistical Society in Significance (Matthews, 2021).
 Amid the outpouring of work that followed, several contributions proposed methods for data analysis that leverage interval estimates for some scalar target of inference $\theta$ to inform decision making in Bayesian or frequentist settings (Blume et al., 2019; Goodman et al., 2019; Matthews, 2019).
@@
 In a Bayesian or frequentist setting, the LPC could be applied to select the smallest sample size $n$ such that $\Pr(L_{1-\alpha}\leq l)\geq q$ for some target length $l$ and probability $q\in(0,1)$.
-This focus on controlling the length of CIs is not to the exclusion of the frequentist notion of coverage, which should approach the nominal level of $1-\alpha$ in frequentist and Bayesian contexts as $n\rightarrow\infty$ under standard regularity conditions (van der Vaart, 1998).
-For Bayesian studies, Stevens and Hagar (2022) recently assigned a distribution to the LPC with respect to the sample size $n$ under the conditional approach.
-This distribution was called the sufficient sample size distribution (SSSD), and it was formally defined so that its $q$-quantile is the smallest sample size $n$ such that $L_{1-\alpha}$ will have length at most $l$ with probability at least $q$ (i.e., such that $\Pr(L_{1-\alpha}\leq l)\geq q$).
-The SSSD was observed to be approximately normal in a variety of situations via simulation (Stevens and Hagar, 2022), but theoretical justification for this approximate normality was not explored.
-Furthermore, conditions under which the SSSD exists and this approximate normality holds were not clearly defined.
-In this work, we define the length probability distribution (abbreviated as the LP distribution) that extends the SSSD to accommodate Bayesian and frequentist CIs under both the conditional and predictive approaches to specify sampling distributions.
+This focus on controlling the length of CIs is not to the exclusion of the frequentist notion of coverage and its Bayesian analogues (Gustafson, 2012), which should approach the nominal level of $1-\alpha$ as $n\rightarrow\infty$ under standard regularity conditions (van der Vaart, 1998).
+In this work, we define a distribution for the LPC with respect to the sample size $n$ and explore its theoretical properties.
+We call this distribution the length probability distribution (abbreivated as the LP distribution).
+The LP distribution is formally defined so that its $q$-quantile is the smallest sample size $n$ such that $L_{1-\alpha}$ will have length at most $l$ with probability at least $q$ (i.e., such that $\Pr(L_{1-\alpha}\leq l)\geq q$).
+This construct broadly accommodates Bayesian and frequentist CIs under both the conditional and predictive approaches to specify sampling distributions.
 By estimating the LP distribution and its quantiles, practitioners can select sample sizes for scientific studies that attain precision criteria.
@@
 The second method leverages this asymptotic theory to inform an efficient method to estimate the LP distribution based on simulation results conducted at only two sample sizes; this method is generally applicable and flexibly accommodates moderate sample size settings where asymptotic approximations may be unsuitable.
-Similar approaches have recently been proposed to efficiently select sample sizes within the hypothesis testing framework (Hagar and Stevens, 2024, 2025a, 2025c, 2025b).
 While our design framework proposed here is theoretically intricate, its implementation is straightforward, promoting an economical and broadly useful approach to simulation-based sample size determination with precision criteria.
@@
 We henceforth denote the length of the $100\times(1-\alpha)\%$ CI as $L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}})$ for a future sample of $n$ ISUs; the length of the corresponding CI for an observed sample is $L_{1-\alpha}(\boldsymbol{w}^{{}_{(n)}})$.
-This notation now emphasizes that the CI length and its sampling distribution are functions of the data.
+This notation now emphasizes that the CI length and its sampling distribution are functions of the data and the confidence or credibility level $1-\alpha$.
 Algorithm 1 details a simulation-based procedure to estimate the sampling distribution of CI lengths at a given sample size $n$.
@@
 This flexibility accommodates the use of simple or complex interval estimates – such as approximate frequentist confidence intervals obtained using normal theory and sample summary statistics or Bayesian highest posterior density intervals based on computationally approximating the posterior distribution of $\theta$.
+While design methodology for likelihood-based inference is the focus of this paper, we discuss extensions for other interval estimates in Section 7.
 The probability of satisfying the length criterion can be estimated as
@@
 Theorem 1 imposes constraints on $\Lambda(\boldsymbol{\eta}_{r})$ and $\tau^{2}$ to ensure the normal distribution that approximates the LP distribution for a given length $l$ has finite mean $\mu_{l}>0$ and standard deviation $\sigma_{l}>0$.
+These constraints impose implicit assumptions that the partial derivatives in (3) and (5) exist.
 For large sample sizes $n$ corresponding to small target lengths $l$, the values for $\mu_{l}$ and $\sigma_{l}$ from Theorem 1 can be used to determine quantiles of the LP distribution under the conditional approach as stated in Corollary 1.
@@
 These proxies substantiate the theory that underpins our proposed methodology, but they are not used in our design methods proposed in Section 5.
+In the hypothesis testing framework, proxies to sampling distributions of test statistics and posterior probabilities have also been used to expedite simulation-based design (Hagar and Stevens, 2024, 2025a, 2025c, 2025b).
 The delta method and the result in (6) give rise to an approximate sampling distribution for $\kappa(\boldsymbol{W}^{{}_{(n)}})=\log(L_{1-\alpha}(\boldsymbol{W}^{{}_{(n)}}))$.
@@
 In previous clinical trials (Klein, 2024), Mpox illness duration was defined as the time in days between starting tecovirimat (a safe drug approved for use with smallpox) and experiencing the resolution of all Mpox lesions.
-For this example, we want to estimate the 0.9-quantile of this distribution for Mpox illness duration.
+For this example, we want to estimate the median of this distribution for Mpox illness duration.
 Our subsequent numerical results serve purely illustrative purposes, but we emphasize the value in precisely estimating quantities associated with illness durations to inform public health recommendations.
@@
 Using our notation from Section 2, the binary indicator $\delta_{i}$ could also be viewed as an additional covariate.
-For illustration, we suppose that the resolution time $T_{i}$ and censoring time $C_{i}$ are exponentially distributed such that the exponential rates are log-linear functions of $x_{i}$:
-$T_{i}\sim\text{EXP}(\gamma+\beta x_{i})~~~\text{and}~~~C_{i}\sim\text{EXP}(\xi+\rho x_{i}),$ (13)
-where $\gamma$, $\beta$, $\xi$, and $\rho$ are regression parameters.
+We model the resolution time $T_{i}$ for participant $i$ using a Bayesian proportional hazards model.
+The maximum follow-up time is 4 weeks, so all participants have an administrative censoring time of $C_{i}=28$ days.
+The baseline “hazard” of experiencing lesion resolution is a piecewise constant function.
+This baseline hazard $h_{0}(t)$ changes with the week of observation such that $h_{0}(t)$ is $h_{1}$ if $0\leq t\leq 7$, $h_{2}$ if $7<t\leq 14$, $h_{3}$ if $14<t\leq 21$, and $h_{4}$ if $t>21$.
+We collectively refer to these constants as $\boldsymbol{h}=(h_{1},h_{2},h_{3},h_{4})$.
+Participant $i$’s hazard of experiencing lesion resolution in week $j$ is therefore $h_{j}\exp(\beta x_{i})$, where $\beta$ is a regression parameter.
 The to-be-observed time $Y_{i}$ for participant $i$ is the minimum of $T_{i}$ and $C_{i}$.
-Because the censoring process in (13) depends only on the observed covariate $x_{i}$, these censored data are considered to be missing at random.
-However, the Bayesian methods we use in this subsection could in theory accommodate more complex missingness mechanisms.
 At the design stage, we must also define a distribution for $\{X_{i}\}_{i=1}^{n}$.
 We suppose that $X\sim\mathcal{N}(0,\sigma_{x}^{2})$ for this example, and that the data from participants $i$ and $i^{\prime}$ are independent for $i\neq i^{\prime}$.
-To summarize, the parameters for our analysis model are $\boldsymbol{\eta}=(\gamma,\beta)$, but the parameters $\boldsymbol{\eta}^{+}=(\gamma,\beta,\xi,\rho,\sigma_{x})$ are used for data generation.
-Our target of inference used for study design is the 0.9-quantile of the distribution for the resolution time $T$ when $x=0$: $\theta=g(\boldsymbol{\eta})=\log(10)\exp(-\mu)$.
-We aim to estimate a $100\times(1-\alpha)\%=99\%$ posterior highest density interval (HDI) for $\theta$ having length at most $l=5$ days with probability $q=0.8$.
+To summarize, the parameters for our analysis model are $\boldsymbol{\eta}=(\boldsymbol{h},\beta)$, but the parameters $\boldsymbol{\eta}^{+}=(\boldsymbol{h},\beta,\sigma_{x})$ are used for data generation.
+Our target of inference $\theta$ used for study design is the median of the distribution for the resolution time $T$ when $x=0$.
+This median is an implicit function $g(\cdot)$ of the baseline hazards $\boldsymbol{h}$; it is the time at which the cumulative baseline hazard equals $\log(2)$.
+We aim to estimate a $100\times(1-\alpha)\%=95\%$ posterior highest density interval (HDI) for $\theta$ having length at most $l=2$ days with probability $q=0.8$.
 To demonstrate the broad applicability of our methods, we design this study under the conditional and predictive approaches.
-For the conditional approach, we use the parameter values $\boldsymbol{\eta}^{+}_{r}=(\log(\log(10)/18),0.5,-\log(32),1,0.25)$ informed by findings from Kim et al.
+For the conditional approach, we use the parameter values $\boldsymbol{\eta}^{+}_{r}=((0.045,0.095,0.0325,0.01),0.5,0.25)$ informed by findings from Kim et al.
 (2023) for all simulation repetitions $r$.
-These parameter values are such that the true value of $\theta_{r}$ is 18 days and lesion resolution times are censored for roughly 20% of participants.
-Under the predictive approach, we assume that $\gamma_{r}\sim\mathcal{N}(\log(\log(10)/18),0.07^{2})$, $\beta_{r}\sim\mathcal{N}(0.5,0.07^{2})$, $\xi_{r}\sim\mathcal{N}(-\log(32),0.07^{2})$, and $\rho_{r}\sim\mathcal{N}(1,0.07^{2})$; we note that small changes to these parameter values give rise to substantial changes in $\theta_{r}$ and the censoring proportions.
+These parameter values are such that the true value of $\theta_{r}$ is 11 days and lesion resolution times are censored for roughly 28% of participants.
+Under the predictive approach, we assume that $\log(\boldsymbol{h}_{r})\sim\mathcal{N}(\log((0.045,0.095,0.0325,0.01)),0.07^{2}\times\mathbb{I}_{4})$ such that $\mathbb{I}_{4}$ is the $4\times 4$ identity matrix and $\beta_{r}\sim\mathcal{N}(0.5,0.07^{2})$; we note that small changes to these parameter values give rise to substantial changes in $\theta_{r}$ and the censoring proportions.
 We again assume that $\sigma_{x,r}=0.25$ for all repetitions under the predictive approach.
-To facilitate Bayesian analysis, we consider a diffuse and informative prior for $\gamma$ to explore how our methods perform.
-The diffuse prior for $\gamma$ is $\mathcal{N}(0,10^{2})$, and the informative prior is $\mathcal{N}(-2,0.3^{2})$.
-This informative prior is relatively well specified for illustration.
-For all simulations in this subsection, we assign a diffuse $\mathcal{N}(0,10^{2})$ prior to $\beta$.
+To facilitate Bayesian analysis, we consider a diffuse and informative prior for $\boldsymbol{\eta}$ to explore how our methods perform.
+The diffuse prior is $\mathcal{N}(0,10^{2}\times\mathbb{I}_{4})$ for $\log(\boldsymbol{h})$ and $\mathcal{N}(0,10^{2})$ for $\beta$.
+The informative prior is $\mathcal{N}(\log((0.045,0.095,0.0325,0.01)),0.3^{2}\times\mathbb{I}_{4})$ for $\log(\boldsymbol{h})$ and $\mathcal{N}(0.5,0.3^{2})$ for $\beta$.
+This joint informative prior for $\log(\boldsymbol{h})$ and $\beta$ is relatively well specified for illustration.
 These choices result in four approach and prior combinations.
@@
 The resulting LP curve approximations are visualized in Figure 2.
-Figure 2: LP curves obtained using various estimation methods when the 0.9-quantile of Mpox illness duration is the target of inference. The horizontal dotted line represents $q=0.8$.
+Figure 2: LP curves obtained using various estimation methods when the median of Mpox illness duration is the target of inference. The horizontal dotted line represents $q=0.8$.
 Algorithm 2 again yields more accurate LP curve approximation than the analytical approximation.
-The performance of the analytical approximation resulting from Theorem 1 is reasonable for the settings with the diffuse priors for $\gamma$.
-Because the analytical approximation to the LP curve does not account for the prior distribution, it overestimates the required sample size when the relatively well-specified informative prior for $\gamma$ is used ($n=478$ vs.
-456 for the conditional approach and $n=506$ vs.
+The performance of the analytical approximation resulting from Theorem 1 is serviceable for the settings with the diffuse prior.
+Because the analytical approximation to the LP curve does not account for the prior distribution, it overestimates the required sample size when the relatively well-specified informative prior is used ($n=392$ vs.
+357 for the conditional approach and $n=413$ vs.
 376 for the predictive approach).
@@
 This greater dispersion in the LP distribution is a direct result of there being greater dispersion in the sampling distribution of CI lengths under the predictive approach.
-It took less than one second to approximate the LP distribution under the conditional approach using Theorem 1.
-We needed to numerically compute 1000 $\lambda(\boldsymbol{\eta}_{r})$ values and their corresponding derivatives in $\tau^{2}$ to approximate the LP distribution under the predictive approach, which took 15 seconds on a high computing server with 72 cores.
-Roughly one hour on this server was required to approximate all LP curves in Figure 2 via Algorithm 2.
+It took one second to approximate the LP distribution under the conditional approach using Theorem 1.
+The analytical approximation for this example is not instantaneous because we needed to numerically approximate the derivatives and Fisher information in (3) and (5).
+When approximating the Fisher information numerically, we can accommodate the straightforward censoring process in this example or more complex censoring mechanisms.
+We numerically computed 1000 $\lambda(\boldsymbol{\eta}_{r})$ values and their corresponding derivatives in $\tau^{2}$ to approximate the LP distribution under the predictive approach, which took 20 seconds on a high computing server with 72 cores.
+Roughly 14 minutes on this server was required to approximate all LP curves in Figure 2 via Algorithm 2.
 This process was computationally intensive because we approximated each posterior of $\boldsymbol{\eta}$ using Markov chain Monte Carlo with 1000 burnin iterations and 5000 retained draws, but it could have been expedited using analytical methods for posterior approximation.
 Nevertheless, Algorithm 2 is still much more efficient than naive simulation.
-On the same computing server, it took 9 and 19 hours to naively approximate the LP curve for this example under the conditional and predictive approaches, respectively.
+On the same computing server, it took 3.5 and 5.5 hours to naively approximate the LP curve for this example under the conditional and predictive approaches, respectively.
 Our numerical results in Appendix C confirm the strong performance of Algorithm 2 in Bayesian settings and further demonstrate that analytical approximations based on Theorem 1 may perform poorly with moderate sample sizes when (i) informative priors are used or (ii) the posterior distributions used to construct the CI exhibit substantial skewness.
@@
 Our proposed framework can be broadly applied to design studies with precision criteria when CIs are constructed using a method that is based on likelihood functions.
-Future work could extend this methodology to accommodate design with CIs that are based on the frequentist nonparametric or the Bayesian bootstrap.
+Nevertheless, similar theory and methods can be developed for certain types of interval estimates that are not likelihood based.
+For instance, suppose the target of inference $\theta$ is a function of the regression parameters $\boldsymbol{\eta}$ from a generalized estimating equation (GEE).
+The variance of the sampling distribution of the M-estimator $\hat{\theta}^{{}_{(n)}}_{r}$ takes the form $n^{-1}\Lambda(\boldsymbol{\eta}_{r})$ (Liang and Zeger, 1986), where $n$ is the number of ISUs and $\Lambda(\boldsymbol{\eta}_{r})$ is based on an ISU-level sandwich variance estimator instead of $\mathcal{I}_{\emph{ISU}}(\boldsymbol{\eta}\mid{\boldsymbol{x}})$.
+A similar result to Theorem 1 can therefore be developed for use with GEEs such that $\mu_{l}$ and $\sigma_{l}$ are functions of the sandwich variance instead of the Fisher information.
+Because Algorithm 2 requires empirical estimation of CI lengths, it can be implemented as is with interval estimates for GEE regression parameters.
+Future work could extend this methodology to accommodate semiparametric and nonparametric interval estimates more broadly, including design with CIs that are based on the frequentist nonparametric or the Bayesian bootstrap.
+It would be important to formally define settings under which analogues to Theorem 1 are theoretically valid to substantiate the use of our methods.
+While semiparametric and nonparametric methods can allow for more flexibility in data analysis, we emphasize that we still need to fully specify data generation processes to implement simulation-based design in these settings.
 Our approach could also be extended in other aspects to accommodate more complex design criteria.
-For instance, it would be useful to consider economical sample size determination methods for studies that aim to satisfy criteria for both the study power and the precision of interval estimates.
-Additionally, we could further develop design methods with precision criteria in a sequential framework or in a framework with multiple targets of inference.
-Such extensions would require additional theoretical development because we would need to account for the level of dependence in the joint sampling distribution of CI lengths across all potential analyses and targets of inference.
-Finally, our design framework requires that the regularity conditions for the asymptotic normality of the MLE and BvM theorem are satisfied.
-It would therefore be of interest to consider design methods with precision criteria in settings where these regularity conditions are relaxed.
+For example, it would be useful to consider economical sample size determination methods for studies that aim to satisfy criteria for both the study power and the precision of interval estimates.
 Supplementary Material